# A substellar flyby that shaped the orbits of the giant planets

### v1

@@ A.1 Likelihood of Multiple Flybys
We can model the likelihood of the solar system experiencing a specific number of flyby encounters using the Poisson distribution. This expresses the probability of a given number of events in a fixed time interval. We can use the Poisson distribution because the number of flyby events does not depend on any previous flyby events and we model the encounter rate using a constant (static) object density, velocity dispersion, and cross-section. Given the encounter rate of $\Gamma\approx 10^{-3}\,\mathrm{flybys}\,\mathrm{Myr}^{-1}$ (see above), the probability of $k$ flyby events in time interval $\Delta t$ is given by ($k\in\mathbb{N},t\in\mathbb{R}^{+}$)
@@
Therefore, the likelihood of zero flybys in $100\,\mathrm{Myr}$ is $P(0,100\,\mathrm{Myr})\approx 0.881$. Similarly, $P(1,100\,\mathrm{Myr})\approx 0.112$, $P(2,100\,\mathrm{Myr})\approx 0.007$, and so on. We can express the probability of having at least $k$ flybys as
@@
such that the likelihood of the solar system experiencing at least one flyby is $\hat{P}(k\geq 1,100\,\mathrm{Myr})\approx 0.119$.
@@ A.2 Success Criteria
We discuss here the success rate for the four criteria defined in Section 3 of (Nesvornỳ and Morbidelli, 2012) resulting from the flyby simulations of this study. Criterion A is that four planets remain in the simulation. Criterion B has three parts: all of the semi-major axes are within $20\%$ of the current solar system, the average eccentricities are all less than 0.11, and the average inclinations are less than 2. See Table 2 for the average eccentricity and inclination values for the best matching flyby case. Criterion C, discussed in this study, requires the amplitude of the $g_{5}$ mode in Jupiter’s orbit to be greater than $0.022$. Criterion D is not applicable to our study because the giant planets do not migrate in our simulations. Instead, a specific study of the effect of each flyby on the terrestrial planets should be considered. The criteria are dependent; C can only be successful if B is successful, and B can only be successful if A is successful. Using these criteria with our data we find success rates of A (41.33%), B (19.33%), and C (1.63%). These success rates also show that $1\%$ of flyby simulations result in a ‘close match’ with the solar system. If we consider an intersection of successful matches from these criteria with $\mathcal{M}$ the success rate is $0.23\%$.
A.3 Angular Momentum Deficit
As discussed in Section 2.3, we found that, for purpose of comparing the similarity between two systems, the angular momentum deficit (AMD) is not a good parameter because the AMD can be dominated by a single planet which is a common result of flyby encounters. There are many more flyby systems with AMD similar to the solar system than there are closely matching flyby systems, see Figure 5. Flyby encounters tend to increase the eccentricities of the outermost planets more than the innermost planets (Heggie and Rasio, 1996). The results from our simulations are generally consistent with this theoretical expectation. Note that this is opposite to the solar system where Neptune has the smallest eccentricity out of the giant planets. Approximately $8\%$ of the closely matching cases have the same trend as in the solar system.
Figure 5: Here we show in black the distribution of the normalized angular momentum deficit (AMD) compared to the solar system (shown in red). We highlight in blue the 422 flyby systems that are a close match to the solar system according to our metric $\mathcal{M}$ along with the best matching system. We note that of the $5\cdot 10^{4}$ flyby systems, $15737$ have AMD within the range of our closely matching systems; $6679$ are within the range of the best matching system; and $567$ are within $1\%$ of the solar system.
Table 2: Averaged over 20 Myr, we show a comparison of the eccentricities and inclinations (with respect to the invariant plane) of the giant planets (SS) and our best matching flyby system that experienced a $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$ encounter at $q_{\star}=1.69\,\mathrm{AU}$, with $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$.

### v2

@@ A.1 Likelihood of Multiple Flybys
We can model the likelihood of the solar system experiencing a specific number of flyby encounters using the Poisson distribution. This expresses the probability of a given number of events in a fixed time interval. We can use the Poisson distribution because the number of flyby events does not depend on any previous flyby events and we model the encounter rate using a constant (static) object density, velocity dispersion, and cross-section. Given the encounter rate of $\Gamma\approx 1.3\cdot 10^{-4}\,\mathrm{Myr}^{-1}$ (see Eq. 5), the probability of $k$ flyby events in time interval $\Delta t$ is given by ($k\in\mathbb{N},t\in\mathbb{R}^{+}$)
@@
Therefore, the likelihood of zero flybys in $100\,\mathrm{Myr}$ is $P(0,100\,\mathrm{Myr})\approx 0.9871$. Similarly, $P(1,100\,\mathrm{Myr})\approx 0.0128$, and so on. We can express the probability of having at least $k$ flybys as
@@
such that the likelihood of the solar system experiencing at least one flyby is $\hat{P}(k\geq 1,100\,\mathrm{Myr})\approx 0.0129$.
@@ A.2 Success Criteria
We discuss here the success rate for the four criteria defined in Section 3 of (Nesvornỳ and Morbidelli, 2012) resulting from the flyby simulations of this study. Criterion A is that four planets remain in the simulation. Criterion B has three parts: all of the semi-major axes are within $20\%$ of the current solar system, the average eccentricities are all less than 0.11, and the average inclinations are less than 2. See Table A.1 for the average eccentricity and inclination values for the best matching flyby case. Criterion C, discussed in this study, requires the amplitude of the $g_{5}$ mode in Jupiter’s orbit to be greater than $0.022$, greater than half of the observed value. Criterion D is not applicable to our study because the giant planets do not migrate in our simulations. Instead, a specific study of the effect of each flyby on the terrestrial planets should be considered. The criteria are dependent; C can only be successful if B is successful, and B can only be successful if A is successful. Using these criteria with our data we find success rates of A (41.33%), B (19.33%), and C (1.63%). These success rates also show that $1\%$ of flyby simulations result in a ‘close match’ with the solar system. If we consider an intersection of successful matches from these criteria with $\mathcal{M}$ the success rate is $0.23\%$.
Table A.1: Averaged over 20 Myr, we show a comparison of the eccentricities and inclinations (with respect to the invariant plane) of the giant planets ($S$) and our best matching flyby system described in section 3.2.
@@
A.3 Angular Momentum Deficit
We noted in Section 2.3 that some previous studies used the angular momentum deficit (AMD) to compare the orbital architectures of planetary systems. However, we found that, for the purpose of comparing the similarity between flyby systems and the solar system, the AMD is not a useful parameter because the AMD can be dominated by a single planet which is a common result of flyby encounters. There are many more flyby systems with AMD similar to the solar system than there are closely matching flyby systems, see Figure A.1. Flyby encounters tend to increase the eccentricities of the outermost planets more than the innermost planets (Heggie and Rasio, 1996). The results from our simulations are generally consistent with this theoretical expectation. Note that this is opposite to the solar system where Neptune has the smallest eccentricity out of the giant planets. Approximately $8\%$ of the closely matching cases have the same trend as in the solar system.
Figure A.1: Here we show in black the distribution of the normalized angular momentum deficit (AMD) compared to the solar system (shown in red). We highlight in blue the 422 flyby systems that are a close match to the solar system according to our metric $\mathcal{M}$ along with the best matching system. We note that of the $5\cdot 10^{4}$ flyby systems, $15737$ have AMD within the range of our closely matching systems; $6679$ are within the range of the best matching system; and $567$ are within $1\%$ of the solar system.
A.4 Orbital elements
We have argued in Section 2.3 that our new metric is well suited for finding close matches of flyby systems to the solar system. Here we show how this metric compares to the approach of matching orbital elements adopted in some previous studies. The semi-major axis, eccentricity, and inclination of the solar system and the flyby simulations are shown in Fig. A.2. We highlight (in blue dots) those flyby simulations that we have labeled a close match using our log-spectral distance metric. For the solar system, we use the ensemble of 20 Myr snapshots from the 5 Gyr solar system integrations described in Sec. 2.4. Note that the orbital elements of the solar system, especially the eccentricities, vary significantly (indicated by their standard deviation over the solar system reference set).
We can see that, compared to the solar system, there is a larger range of orbital parameters in the flyby outcomes. The close match cases do overlap with the solar system in the space of orbital parameters. For Jupiter, Saturn, and Uranus, the median eccentricities and inclinations of the close match cases are comparable to those of the actual solar system. Neptune has a higher eccentricity in many (but not all) flyby simulations compared to the actual solar system because it is the outermost planet and thus most affected by distant flybys which tend to be the most common ones.
Figure A.2: The distributions of the osculating eccentricity, inclination, and semi-major axis of the outer solar system planets in each flyby simulation (black crosses) as well as the subset of those that led to a close match with the solar system (blue dots, see also Fig. 2). For the solar system, we plot the average and standard deviation of the initial osculating orbital elements of the reference set of solar systems described in Section 2.4.
We stress that the reverse is not true: a good match in the space of orbital elements is not necessarily a good match when evaluated using our log-spectral metric. To illustrate this point, we find the closest match by only comparing orbital elements of the flyby simulations to the orbital elements of the solar system (semi-major axis, eccentricity, and inclination). We do this using a least squares approach to define the following metric to compare the orbital elements:
$\mathcal{M}_{\rm oe}$ $=$ $\sum_{k=5}^{8}\left(\frac{e_{k}-\bar{e}_{k}}{\bar{e}_{k}}\right)^{2}+\sum_{k=5}^{8}\left(\frac{i_{k}-\bar{i}_{k}}{\bar{i}_{k}}\right)^{2}$ (8)
$+0.1\sum_{i=5}^{8}\left(\frac{a_{i}-\bar{a}_{i}}{\bar{a}_{i}}\right)^{2},$
where the sum is over the four outer planets and the barred parameters are those of the present day outer solar system planets. Note that this metric is by no means unique and in fact quite arbitrary. For example, we added a relative weight of 0.1 to the fractional semi-major axis to avoid overfitting this component.
We plot the spectrum of the closest match that we find using the $\mathcal{M}_{\rm oe}$ metric in Fig. A.3. The flyby parameters for this encounter are $m_{\star}=1.67\,\mathrm{M}_{\mathrm{J}}$, $q_{\star}=4.7\,\mathrm{AU}$, $v_{\infty}=2.4\,\mathrm{km\,s}^{-1}$, and $i_{\star}=52^{\circ}$. One can see that both the location and the amplitude of the peaks are indeed a good match to the solar system since they are determined by the orbital parameters we compare. However, the spectrum looks otherwise quite different than that of the solar system. There is a wide range of other frequencies with amplitudes that are several orders of magnitude higher than those of the solar system. One can also see frequencies and their harmonics which are not present in the solar system. This example demonstrates the need to go beyond comparing orbital parameters or comparing only the fundamental secular modes. We think that our new log-spectral distance metric can more reliably identify a simulated system that matches the secular architecture of the solar system.
Figure A.3: Power spectra of the complex eccentricities (left column) and complex inclinations (right column). The black line shows the best matching flyby case found by comparing only orbital parameters (see Eq. 8). The color-shaded plots show the ensemble of spectra of the solar system integrations. Only a subset of the peaks of the flyby system’s power spectra match those of the solar system. Aside from those matching peaks, the spectra are qualitatively very different from each other. Our log-spectral metric is able to find much better matches as shown in Fig. 5.
A.5 Null Hypothesis
Here we test the “null hypothesis", namely, given our circular, coplanar initial conditions, can the effect of planet-planet interactions alone (without any flybys) generate the secular mode amplitudes found in the observed solar system? We start with a random subset of $1000$ configurations from the same initial conditions we used in our set of $5\cdot 10^{4}$ flyby simulations. However, because a perfectly coplanar, point-mass planetary system simulation will never generate any relative inclinations we slightly perturb the planets’ initial vertical positions and velocities on the order of $10^{-8}$ fraction of their orbital radius and orbital velocity, respectively. As with other simulations throughout this work, we integrate this set for $20\,\mathrm{Myr}$ taking $2^{11}$ evenly-spaced samples. The power spectra of the resulting complex eccentricities and complex inclinations in this ensemble are shown in Figure A.4. For comparison, the power spectra of the actual solar system today are also plotted in this figure.
We briefly comment on these results. As expected (and noted in Section 3), the choice to initialize the giant planets on their current semi-major axes results in fundamental secular mode frequencies similar to those observed in the solar system today. We see this in Figure A.4 for both the eccentricity and inclination power spectra. However, the maximum amplitudes of the eccentricities are on the order of $10^{-3}$ (roughly $\sim\mathrm{M_{J}}/\mathrm{M}_{\odot}$), and the inclination amplitudes are all below about $10^{-6}$; the latter underscore the low efficiency of inclination excitation from a coplanar initial state. Comparing with the power spectra of the actual solar system, it is evident that the planet-planet interactions of the giant planets alone cannot excite the eccentricity and relative inclination amplitudes observed in the solar system today.
Figure A.4: Testing the "null hypothesis": the power spectra of an ensemble of 1000 integrations of the giant planets initialized in circular and very nearly coplanar orbits, integrated for 20 Myr, with no stellar flybys. The median power at each frequency of the ensemble is shown as a black line, with gray-shaded regions representing the one–, two– and three–ς deviations. For comparison, the color-shaded plots are the actual solar system’s power spectra. This shows that, from initially very nearly circular and coplanar orbits (at their observed semi-major axes), planet-planet interactions alone cannot account for the observed amplitudes of the giant planets’ secular modes. See Section A.5 for details.
Figure A.5: Here we show the distributions of all the flyby parameters in our simulations (black) as well as the subset of those parameters that led to a close match with the solar system (blue). The contours show the first two standard deviations. See Section 3 for more details.